OpenStudy (anonymous):

a motor cycle travels 550 miles in the same time that a truck travels 500 miles. If the motorcycle's speed is 5 miles per hour faster than the truck's speed, find the motorcycle's speed and the trucks speed

OpenStudy (anonymous):

Well to start off remember that velocity = distance over time, though for this question solving for time first is the way to go. So the equation is t=d/v and remember that both times are equal! Solve for v :)

OpenStudy (anonymous):

Let x be the speed of the motorcycle.\[\frac{550}{x }= \frac{500}{x-5}\] x = 55 mph x-5 = 50 mph